OpenStudy (anonymous):

Carols has $3 and Ryan has $-4. How much more does Carols have than Ryan?

OpenStudy (phantomcrow):

What do you need to add to -4 to get to 3?

OpenStudy (anonymous):

1 sec

OpenStudy (anonymous):

7?

OpenStudy (phantomcrow):

Yes.

OpenStudy (anonymous):

i just thought that because Ryan doesn't have any money, the answer should be 3 because that is all the money Carlos has

OpenStudy (texaschic101):

How much more does carlos have then ryan.....you would subtract ryan from carlos 3 - (-4) = 3 + 4 = 7

OpenStudy (phantomcrow):

Realistically, yes. But we're talking in a theoretical and hypothetical way.

OpenStudy (anonymous):

yes i understand i was just looking at the realistic point of view

OpenStudy (texaschic101):

I can understand the confusion